Julio left the hospital and drove toward the recycling plant at an average speed of 55 mph. Jasmine left four hours later and drove the opposite direction with an average speed of 65 mph. How long does Jasmine need to drive before they are 340 mi. apart?
Found 2 solutions by solver91311, stanbon:
Answer by solver91311(24713)



By the time Jasmine leaves, Julio has been driving for 4 hours at 55 miles per hour. So when Jasmine starts they are already 220 miles apart. That means together they have to go 340 - 220 = 120 miles farther. Since they are travelling in opposite directions, they are moving apart at 55 plus 65 = 120 miles per hour. So how long does it take to cover 120 miles at 120 miles per hour?

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by stanbon(75887)

Julio left the hospital and drove toward the recycling plant at an average speed of 55 mph. Jasmine left four hours later and drove the opposite direction with an average speed of 65 mph. How long does Jasmine need to drive before they are 340 mi. apart?
---
Julio DATA:
rate = 55 mph ; time = x hrs ; distance = 55x miles
-------------------
Jasmine DATA:
rate = 65 mph ; time = x-4 hrs ; distance = 65(x-4) miles
-----------------------
Equation:
55x + 65(x-4) = 340 miles
55x + 65x - 260 = 340 miles
120x = 600
x = 5 hr
---
Jasmine time = x-4 = 5-4 = 1 hr.
===================================
Cheers,
Stan H.
==============